[labmembers] Litho update
mahnaz at stanford.edu
Fri Jan 11 14:00:07 PST 2013
happy new year and welcome back...
Yes oven went out for service and re-assembled. The contact angle 66.47
Both headway and Laurell got cleaned and reassembled.
All the hotplates were calibrated to the spec.
all the ovens temperature were checked.
all the litho box went out and got cleaned.
all the metal boats but two went out and got cleaned.
all the litho cassettes are cleaned.
I have attached an excel sheet with all the tracks thickness and
exposure times on the KS aligners and EVG aligner
so please take a look at it.
I will put a copy of it next to the aligners.
Asml is finishing up the last part of the annual pm by running the
metrology tests on the tool.
-------------- next part --------------
A non-text attachment was scrubbed...
Name: Copy of Litho_quals sheet1-7-13.xlsx
Size: 24195 bytes
Desc: not available
More information about the labmembers